# Understanding Outer Optimizers in Local SGD: Learning Rates, Momentum, and Acceleration

### v1

@@ 1 Introduction
Training very large scale machine learning models requires a lot of compute. This compute is often centrally controlled by a single entity and tightly connected in a data center. Gradients are constantly synchronized, hardware failures are controlled and mitigated, and things (mostly) run smoothly. Building this training infrastructure is expensive, however, and centralized control might not be desirable for all models. This has led to a surge of interest in decentralized collaborative training of large-scale models across different, potentially poorly connected clusters [10, 23, 22]. This has motivated the adoption of federated learning algorithms in training language models, chiefly for scalability and communication efficiency rather than data privacy. Efficient parallelization strategies also factored in the remarkable recent training of DeepSeek V3 and R1 on a tight budget [7, 6].
A foundational algorithm in distributed and federated optimization is Local SGD [54]. Many popular algorithms fit in the FedOpt template [48] (Algorithm 1), including FedAdam [48], FedRR [40, 38], DiLoCo [10, 23] and many others. FedOpt solves the minimization problem $\min_{x\in\mathbb{R}^{d}}f(x)$ given access to $M$ different computational nodes and unbiased stochastic gradients of $f$. FedOpt consists of three main components: an inner update loop on every client, an aggregation of the client updates, and then an outer update step taken on the server.
@@
When both the local and outer update rules correspond to gradient descent (i.e. $x_{\mathrm{new}}=x_{\mathrm{old}}-\beta\Delta$ for some stepsize $\beta$ and update vector $\Delta$), the corresponding algorithm is Generalized Local SGD. If we additionally take the outer stepsize to be $1$, we get Local SGD. Local SGD simply does $H$ steps of SGD on each node, and then averages the result after applying the updates. This is the most common form in which the algorithm is analyzed, as in e.g. [52, 27, 58, 28, 13, 46]. In practice, different choices of outer optimizers perform better. For example, DiLoCo/OpenDiLoCo use SGD with Nesterov Momentum as the outer optimizer [10]. This has motivated much analysis of different outer optimizers and their impact [48, 37, 24, 53]. However, our theoretical understanding of the fundamental Generalized Local SGD algorithm remains limited. In particular, it is not clear why the bilevel optimization structure of the algorithm is helpful from an optimization perspective, even in the i.i.d. setting where the data distribution is the same on all the nodes. Additionally and to the best of our knowledge, we have no explicit expressions for what the ideal learning rate pair $(\eta,\gamma)$ should be. Empirically, outer optimizers employing Nesterov acceleration have the best performance, yet to the best of our knowledge why or how it improves convergence is not known.
@@
  • We conduct a novel, tighter analysis of Generalized Local SGD (Theorem 3.3) that shows the outer learning rate plays a dual role. It (a) interpolates between two extreme regimes: taking many effective steps at the cost of higher variance to taking fewer steps but at reduced variance and (b) increases the algorithmic robustness to hyperparameter tuning by making up for ill-tuned inner learning rates. The latter holds even in the absence of any stochastic gradient noise.
  • We extend the above analysis to cover Generalized Local SGD where the outer optimizer also uses momentum (Theorem 3.5) and show that this gives additional leeway in tuning $\gamma$.
  • We provide a convergence analysis for Local SGD with an accelerated outer optimizer and unaccelerated inner optimizer (Theorem 3.6), showing that using Nesterov acceleration in the outer loop achieves better dependence on the number of communication rounds $R$ in the drift terms compared to standard Local SGD and improving upon the convergence rate of FedAc [60].
  • We also derive an adaptive, data-dependent, high-probability guarantee for the convergence of Local SGD with GD as the outer optimizer (Theorem 3.8) that shows further benefits of tuning the outer stepsize in more nuanced settings.
@@ 2 Related Work
There is a rich literature on algorithms for communication-efficient distributed optimization for federated learning [29], where multiple clients collaborate on solving a machine learning problem  [54]. Federated learning algorithms are designed to reduce the effect of data heterogeneity [26, 54, 42], ensure the data stays private [56], deal with intermittent or cyclic client availability [12], among other issues.
@@
The most popular algorithm in the federated optimization literature is Local SGD or Federated Averaging [54]. It is a generalization of minibatch SGD that, rather than communicating at every step of the optimization process, communicates only intermittently. Local SGD shows remarkable efficiency in many settings in practice, and therefore its convergence and generalization properties have been the subject of intense theoretical investigation over the past few years [52, 27, 58, 57, 45, 13, 15, 46]. Many variants of Local SGD exist, including those that use random reshuffling instead of i.i.d. sampling locally [61, 40], adaptive methods such as Adam [48, 55], and modifications to handle data heterogeneity [26, 41], personalization [17], or additionally use gradient compression [16, 50]. Generalized Local SGD, where we use two stepsizes (as in Algorithm 1), is known to be important in managing the trade-off between converging quickly and converging to a mismatched point in heterogeneous distributed optimization [57, 5, 46]. Our focus here is on the homogeneous or i.i.d. data setting; Here, the most related works are [26, 37, 24, 53] and we discuss our work’s relation to theirs in detail in the next section after reviewing some preliminaries.
@@ 3.1 Preliminaries
To simplify our analysis, we follow [52] and define the virtual sequences
@@ 3.2 Main convergence result
Existing results on the convergence of Gen. Local SGD. When the outer stepsize $\gamma=1$, the convergence of (GEN-LOC-SGD) is very well understood, with tightly matching upper and lower bounds [27, 58, 13]. In particular, the best rate for the algorithm is
@@
The first two terms in the above convergence guarantee show that increasing the number of local steps has the same effect as increasing the number of communication rounds $R$, and are identical to the convergence guarantee of doing $RH$ steps of SGD with minibatch size $M$. Local SGD differs from ordinary minibatch SGD in the last term, which shows different scaling between $H$ and $R$, where increasing $R$ helps more than increasing $H$. This is because increasing $H$ incurrs additional client drift that slows down the convergence of the algorithm in the presence of stochastic gradient noise. When the outer stepsize $\gamma$ is allowed to vary, the convergence of the algorithm is less clear. [26] gives the following convergence rate in the absence of data heterogeneity,
$\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{R}+\frac{\sigma\left\lVert x_{0}-x_{\ast}\right\rVert}{\sqrt{MR}}\right),$
for specially chosen $\eta$ and $\gamma$ pairs. This rate matches that of Minibatch SGD, but does not recover the convergence rate of vanilla Local SGD given by Equation˜2. [24] also give a guarantee for Generalized Local SGD with a specific outer learning rate that is always at least $1$ and that depends on the heterogeneity of the iterates across the different clients. Since the analysis is conducted in the heterogeneous setting, the local stepsize required to scale with $1/H$. A guarantee that applies to any outer learning rate in the nonconvex, heterogeneous setting given by [53].
@@
Theorem 3.3.
Suppose that Assumptions˜3.1 and 3.2 hold. Then the iterates generated by Generalized Local SGD run with local stepsize $\eta>0$ and outer stepsize $\gamma>0$ for $R$ communication rounds and with $H$ local steps per round satisfy,
$\!\mathbb{E}\!\left[f\!\left(\!\frac{1}{RH}\!\sum_{r=0}^{R-1}\!\sum_{h=0}^{H-1}\!y_{r,h}\!\right)\!\right]\!-f(x_{\ast})\leq\mathcal{O}\left(\!\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma RH}+\frac{\eta\sigma^{2}(1+(\gamma-1)_{+})}{M}+L\eta^{2}\sigma^{2}H\!\right),$ (3)
where $(a)_{+}=\max(a,0)$. and provided the stepsizes $\eta$ and $\gamma$ jointly satisfy $\eta L(1+(\gamma-1)_{+}H)\leq\frac{1}{4}$.
Implications of Theorem˜3.3. Before giving a proof sketch for Theorem˜3.3, we first discuss its implications. Observe that when $\gamma\leq 1$, we are allowed to choose $\eta$ larger than $\Omega(\frac{1}{LH})$. This is crucial to obtain the rate of Equation˜2. Indeed, when $\gamma=1$, the requirement on $\eta$ reduces to $\eta L\leq\frac{1}{4}$ and we can choose $\eta$ following [58] as
$\eta=\min\left(\frac{1}{4L},\sqrt{\frac{M{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\sigma^{2}RH}},\left[\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{L\sigma^{2}H^{2}R}\right]^{\frac{1}{3}}\right)$
@@
$f(y_{\mathrm{out}})-f(x_{\ast})\leq\frac{8L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma_{\ast}R}+\frac{\sigma^{2}H}{8R^{2}L}+\frac{\gamma_{\ast}\sigma^{2}}{4LMH},$
@@
$f(y_{\mathrm{out}})-f(x_{\ast})\leq\frac{LD^{2}}{R}+\frac{\sigma D}{\sqrt{MRH}}.$
This confirms the intuition that at the extremes, manipulating the stepsizes $\gamma$ and $\eta$ allows us to interpolate between minibatch SGD and (vanilla) Local SGD, as observed by [57]. In fact, Theorem˜3.3 allows us to go a step further and get an explicit expression for the optimal inner and outer stepsizes depending on the problem parameters. This is given by the following proposition.
Proposition 3.4.
@@
$h(\eta,\gamma)=\frac{D^{2}}{\eta\gamma RH}+L\sigma^{2}H\eta^{2}+\frac{\eta(1+(\gamma-1)_{+})\sigma^{2}}{M}.$ (4)
@@
$\min_{\eta>0,\gamma>0}h(\eta,\gamma)\qquad\text{subject to}\quad$ $\eta L\left(1+(\gamma-1)_{+}H\right)\leq\frac{1}{4}.$ (5)
@@
    $2LH\sigma^{2}\eta^{3}+\frac{\sigma^{2}}{M}\eta^{2}-\frac{D^{2}}{RH}=0.$
@@
    $\gamma_{B}(\eta)=1+\frac{1}{H}\left(\frac{1}{4L\eta}-1\right),$
@@
    $-\frac{4L^{2}D^{2}(H-1)}{R}+2L\sigma^{2}H\eta\left(\eta L(H-1)+1\right)^{2}+\frac{\sigma^{2}(H-1)}{MH}(\eta L(H-1)+1)^{2}=0.$
@@
The proof of the above proposition is straightforward and follows by writing the KKT conditions for the optimization problem in Equation˜5. A consequence of Proposition˜3.4 is that in the case of ill-tuning of the inner stepsize $\eta$, a large outer stepsize $\gamma$ can make up for it. For example, if $\sigma\rightarrow 0$ and $\eta LH\ll\mathcal{O}(1)$, we can make up for this by choosing $\gamma$ as $\frac{1}{\eta LH}$. Thus, we can interpret the outer learning rate $\gamma$ as having two dual roles. (a) It allows us to interpolate between minibatch SGD $(\gamma>1)$ and vanilla Local SGD $(\gamma=1)$, giving us the better of the two rates, and (b) it provides us some additional leeway in hyperparameter tuning by making up for ill-tuned inner learning rate $\eta$.
Our theory suggests that in the worst case, choices of $\gamma<1$ are not useful from an optimization perspective. We should either choose $\gamma=1$ or $\gamma>1$. This can be seen even on quadratic objectives, for example if $f(x)=\frac{x^{\top}Qx}{2}$ for some positive definite matrix $Q$, then a straightforward computation gives the expected iterate after $H$ local steps and $R$ communication rounds is
$\mathbb{E}\left[x_{R}\right]=((1-\gamma)I+\gamma(I-\eta Q)^{H})x_{0}.$
From this, it is clear that if $\eta$ is chosen such that $(I-\eta Q)^{H}$ has eigenvalues smaller than $1$, we should choose $\gamma\geq 1$. While if $(I-\eta Q)^{H}$ has any eigenvalues larger than $1$, we should just choose $\gamma=0$ (i.e. just don’t apply the algorithm at all). In other words, $\gamma$ can make up for a learning rate that is too small, but not a learning rate that is too large. This observation does not exclude that $\gamma<1$ can be useful from a generalization perspective, as noted for the case of a single client by [62], in the presence of data heterogeneity, as noted by [4], or in the presence of time-varying noise (see next subsection).
Proof sketch for Theorem˜3.3. We first start by expanding the update for the round iterate $x_{r+1}-x_{\ast}=x_{r+1}-x_{r}+x_{r}-x_{\ast}$ similar to [26] to get,
@@
where $g_{r,h}$ is defined as in Equation˜1. [26, 24] control the inner product $\left\langle x_{r}-y_{r,h},g_{r,h}\right\rangle$ by either using smoothness or Young’s inequality; This would force us to bound the stray ${\left\lVert y_{r,h}-x_{r}\right\rVert}^{2}$ and take the local stepsize $\eta$ to be small in order to ensure convergence. Instead, we rely on bounding this quantity directly by viewing it as the regret in the online convex optimization sense with respect to the comparator $x_{r}$. Observe that the virtual sequence of averaged local iterates satisfies $y_{r,h+1}=y_{r,h}-\eta g_{r,h}$, and thus through standard regret analysis we have
$\sum_{h=0}^{H-1}-\left\langle x_{r}-y_{r,h},g_{r,h}\right\rangle$ $=\frac{-{\left\lVert y_{r,h}-x_{r}\right\rVert}^{2}}{2\eta}+\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}.$ (6)
The negative terms ${\left\lVert y_{r,H}-x_{r}\right\rVert}^{2}$ in Equation˜6 turn out to be crucial in obtaining an analysis that works for all $\eta$ and not just small $\eta$. With this change and through carefully bounding the variance terms following [27, 58], we obtain the guarantee of Theorem˜3.3. The full proof is provided in Section˜B.2.
Comparison with results on related algorithms. [37] analyze a closely related variant of the algorithm that uses federated random reshuffling [40] as a base. This is a significantly different algorithm that doesn’t allow for an arbitrary number of local steps $H$ and depends on $f$ posessing finite-sum structure. Nevertheless, we can still specialize [38, Theorem 2] approximately to our setting, by using $H$ as the number of data points in an epoch. In our notation, their convergence guarantee reads
$\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma HR}+\eta^{2}H^{2}\sigma^{2}\right),$
@@
$\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{R}\right),$
this rate is the same as gradient descent for $R$ steps (since the finite-sum structure means that per-epoch we approximate one step of gradient descent when $\eta$ is small). A similar rate is derived in [30, 31] if we have access to the proximal operator (i.e. we can do many local steps $H$ on a modified objective). [30] in particular show that an outer learning rate greater than $1$ can be particularly useful for improving the convergence of FedProx [32] in the heterogeneous setting when the smoothness constant varies significantly between different clients.

### v2

@@ 1 Introduction
Training very large scale machine learning models requires a lot of compute. This compute is often centrally controlled by a single entity and tightly connected in a data center. Gradients are constantly synchronized, hardware failures are controlled and mitigated, and things (mostly) run smoothly. Building this training infrastructure is expensive, however, and centralized control might not be desirable for all models. This has led to a surge of interest in decentralized collaborative training of large-scale models across different, potentially poorly connected clusters [douillard23_diloc, jaghouar24_opend, jaghouar2024intellect]. This has motivated the adoption of federated learning algorithms in training language models, chiefly for scalability and communication efficiency rather than data privacy. Efficient parallelization strategies also factored in the remarkable recent training of DeepSeek V3 and R1 on a tight budget [liu2024deepseek, guo2025deepseek].
A foundational algorithm in distributed and federated optimization is Local SGD [wang21_field_guide_to_feder_optim]. Many popular algorithms fit in the FedOpt template [reddi20_adapt_feder_optim] (Algorithm 1), including FedAdam [reddi20_adapt_feder_optim], FedRR [mishchenko21_proxim_feder_random_reshuf, malinovsky22_feder_random_reshuf_with_compr_varian_reduc], DiLoCo [douillard23_diloc, jaghouar24_opend] and many others. FedOpt solves the minimization problem $\min_{x\in\mathbb{R}^{d}}f(x)$ given access to $M$ different computational nodes and unbiased stochastic gradients of $f$. FedOpt consists of three main components: an inner update loop on every client, an aggregation of the client updates, and then an outer update step taken on the server.
@@
When both the local and outer update rules correspond to gradient descent (i.e. $x_{\mathrm{new}}=x_{\mathrm{old}}-\beta\Delta$ for some stepsize $\beta$ and update vector $\Delta$), the corresponding algorithm is Generalized Local SGD. If we additionally take the outer stepsize to be $1$, we get Local SGD. Local SGD simply does $H$ steps of SGD on each node, and then averages the result after applying the updates. This is the most common form in which the algorithm is analyzed, as in e.g. [stich18_local_sgd_conver_fast_commun_littl, khaled19_tight_theor_local_sgd_ident_heter_data, woodworth20_is_local_sgd_better_than_minib_sgd, koloskova20_unified_theor_decen_sgd_with, glasgow21_sharp_bound_feder_averag_local, patel24_limit_poten_local_sgd_distr]. In practice, different choices of outer optimizers perform better. For example, DiLoCo/OpenDiLoCo use SGD with Nesterov Momentum as the outer optimizer [douillard23_diloc]. This has motivated much analysis of different outer optimizers and their impact [reddi20_adapt_feder_optim, malinovsky22_server_side_steps_sampl_without, jhunjhunwala23_fedex, sun23_role_server_momen_feder_learn]. However, our theoretical understanding of the fundamental Generalized Local SGD algorithm remains limited. In particular, it is not clear why the bilevel optimization structure of the algorithm is helpful from an optimization perspective, even in the i.i.d. setting where the data distribution is the same on all the nodes. Additionally and to the best of our knowledge, we have no explicit expressions for what the ideal learning rate pair $(\eta,\gamma)$ for the inner and outer updates, respectively, should be. Empirically, outer optimizers employing Nesterov acceleration have the best performance, yet to the best of our knowledge why or how it improves convergence is not known.
@@
  • We conduct a novel, tighter analysis of Generalized Local SGD (Theorem 1) that shows the outer learning rate plays a dual role. It (a) interpolates between two extreme regimes: taking many effective steps at the cost of higher variance to taking fewer steps but at reduced variance and (b) increases the algorithmic robustness to hyperparameter tuning by making up for ill-tuned inner learning rates. The latter holds even in the absence of any stochastic gradient noise.
  • We extend the above analysis to cover Generalized Local SGD where the outer optimizer also uses momentum (Theorem 2) and show that this gives additional leeway in tuning $\gamma$.
  • We provide a convergence analysis for Local SGD with an accelerated outer optimizer and unaccelerated inner optimizer (Theorem 3), showing that using Nesterov acceleration in the outer loop achieves better dependence on the number of communication rounds $R$ in the drift terms compared to standard Local SGD and improving upon the convergence rate of FedAc [yuan20_feder_accel_stoch_gradien_descen].
  • We also derive a data-dependent, high-probability guarantee for the convergence of Local SGD with GD as the outer optimizer (Theorem 4) that shows further benefits of tuning the outer stepsize in more nuanced settings.
@@ 2 Related Work
There is a rich literature on algorithms for communication-efficient distributed optimization for federated learning [konecny16_feder_learn], where multiple clients collaborate on solving a machine learning problem  [wang21_field_guide_to_feder_optim]. Federated learning algorithms are designed to reduce the effect of data heterogeneity [karimireddy19_scaff, wang21_field_guide_to_feder_optim, murata21_bias_varian_reduc_local_sgd], ensure the data stays private [wei2020federated], deal with intermittent or cyclic client availability [eichner2019semi], among other issues.
@@
The most popular algorithm in the federated optimization literature is Local SGD or Federated Averaging [wang21_field_guide_to_feder_optim]. It is a generalization of minibatch SGD that, rather than communicating at every step of the optimization process, communicates only intermittently. Local SGD shows remarkable efficiency in many settings in practice, and therefore its convergence and generalization properties have been the subject of intense theoretical investigation over the past few years [stich18_local_sgd_conver_fast_commun_littl, khaled19_tight_theor_local_sgd_ident_heter_data, woodworth20_is_local_sgd_better_than_minib_sgd, woodworth20_minib_vs_local_sgd_heter_distr_learn, patel2023on, glasgow21_sharp_bound_feder_averag_local, gu23_why_when_does_local_sgd, patel24_limit_poten_local_sgd_distr]. Many variants of Local SGD exist, including those that use random reshuffling instead of i.i.d. sampling locally [yun21_minib_vs_local_sgd_with_shuff, mishchenko21_proxim_feder_random_reshuf], adaptive methods such as Adam [reddi20_adapt_feder_optim, wang22_commun_effic_adapt_feder_learn], and modifications to handle data heterogeneity [karimireddy19_scaff, mitra21_achiev_linear_conver_feder_learn], personalization [hanzely20_lower_bound_optim_algor_person_feder_learn], or additionally use gradient compression [haddadpour20_feder_learn_with_compr, safaryan21_smoot_matric_beat_smoot_const]. Generalized Local SGD, where we use two stepsizes (as in Algorithm 1), is known to be important in managing the trade-off between converging quickly and converging to a mismatched point in heterogeneous distributed optimization [woodworth20_minib_vs_local_sgd_heter_distr_learn, charles20_outsiz_impor_learn_rates_local_updat_method, patel24_limit_poten_local_sgd_distr]. Our focus here is on the homogeneous or i.i.d. data setting; Here, the most related works are [karimireddy19_scaff, malinovsky22_server_side_steps_sampl_without, jhunjhunwala23_fedex, sun23_role_server_momen_feder_learn] and we discuss our work’s relation to theirs in detail in the next section after reviewing some preliminaries.
@@ 3.1 Preliminaries
To simplify our analysis, we follow [stich18_local_sgd_conver_fast_commun_littl] and define the virtual sequences
@@ 3.2 Main convergence result
Existing results on the convergence of Gen. Local SGD. When the outer stepsize $\gamma=1$, the convergence of (GEN-LOC-SGD) is very well understood, with tightly matching upper and lower bounds [khaled19_tight_theor_local_sgd_ident_heter_data, woodworth20_is_local_sgd_better_than_minib_sgd, glasgow21_sharp_bound_feder_averag_local]. In particular, the best rate for the algorithm is
@@
The first two terms in the above convergence guarantee show that increasing the number of local steps has the same effect as increasing the number of communication rounds $R$, and are identical to the convergence guarantee of doing $RH$ steps of SGD with minibatch size $M$. Local SGD differs from ordinary minibatch SGD in the last term, which shows different scaling between $H$ and $R$, where increasing $R$ helps more than increasing $H$. This is because increasing $H$ incurrs additional client drift that slows down the convergence of the algorithm in the presence of stochastic gradient noise. When the outer stepsize $\gamma$ is allowed to vary, the convergence of the algorithm is less clear. [karimireddy19_scaff] gives the following convergence rate in the absence of data heterogeneity,
$\textstyle\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{R}+\frac{\sigma\left\lVert x_{0}-x_{\ast}\right\rVert}{\sqrt{MR}}\right),$
for specially chosen $\eta$ and $\gamma$ pairs. This rate matches that of Minibatch SGD, but does not recover the convergence rate of vanilla Local SGD given by Equation˜2. [jhunjhunwala23_fedex] also give a guarantee for Generalized Local SGD with a specific outer learning rate that is always at least $1$ and that depends on the heterogeneity of the iterates across the different clients. Since the analysis is conducted in the heterogeneous setting, the local stepsize required to scale with $1/H$. A guarantee that applies to any outer learning rate in the nonconvex, heterogeneous setting given by [sun23_role_server_momen_feder_learn].
@@
Theorem 1.
Suppose that Assumptions˜1 and 2 hold. Then the iterates generated by Generalized Local SGD run with local stepsize $\eta>0$ and outer stepsize $\gamma>0$ for $R$ communication rounds and with $H$ local steps per round satisfy,
$\textstyle\!\mathbb{E}\!\left[f\!\left(\!\frac{1}{RH}\!\sum_{r=0}^{R-1}\!\sum_{h=0}^{H-1}\!y_{r,h}\!\right)\!\right]\!-f(x_{\ast})\leq\mathcal{O}\left(\!\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma RH}+\frac{\eta\sigma^{2}\max(\gamma,1)}{M}+L\eta^{2}\sigma^{2}H\!\right),$ (3)
provided the stepsizes $\eta$ and $\gamma$ jointly satisfy $\eta L(1+(\gamma-1)_{+}H)\leq\frac{1}{4}$ and where $(a)_{+}=\max(a,0)$.
Implications of Theorem˜1. Before giving a proof sketch for Theorem˜1, we first discuss its implications. Observe the stepsize condition $\eta L(1+(\gamma-1)_{+}H)$ is asymmetric in $\eta$ and $\gamma$; That is, when $\gamma\leq 1$, we are allowed to choose $\eta$ larger than $\Omega(\frac{1}{LH})$. This is crucial to obtain the rate of Equation˜2. Indeed, when $\gamma=1$, the requirement on $\eta$ reduces to $\eta L\leq\frac{1}{4}$ and we can choose $\eta$ following [woodworth20_is_local_sgd_better_than_minib_sgd] as
$\textstyle\eta=\min\left(\frac{1}{4L},\sqrt{\frac{M{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\sigma^{2}RH}},\left[\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{L\sigma^{2}H^{2}R}\right]^{\frac{1}{3}}\right)$
@@
$\textstyle f(y_{\mathrm{out}})-f(x_{\ast})\leq\frac{8L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\gamma_{\ast}R}+\frac{\sigma^{2}H}{8R^{2}L}+\frac{\gamma_{\ast}\sigma^{2}}{4LMH},$
@@
$\textstyle f(y_{\mathrm{out}})-f(x_{\ast})\leq\frac{LD^{2}}{R}+\frac{\sigma D}{\sqrt{MRH}}.$
This confirms the intuition that at the extremes, manipulating the stepsizes $\gamma$ and $\eta$ allows us to interpolate between minibatch SGD and (vanilla) Local SGD, as observed by [woodworth20_minib_vs_local_sgd_heter_distr_learn]. In fact, Theorem˜1 allows us to go a step further and get an explicit expression for the optimal inner and outer stepsizes depending on the problem parameters. This is given by the following proposition.
Proposition 1.
@@
$\textstyle h(\eta,\gamma)=\frac{D^{2}}{\eta\gamma RH}+L\sigma^{2}H\eta^{2}+\frac{\eta\max(\gamma,1)\sigma^{2}}{M}.$ (4)
@@
$\textstyle\min_{\eta>0,\gamma>0}h(\eta,\gamma)\qquad\text{subject to}\quad$ $\eta L\left(1+(\gamma-1)_{+}H\right)\leq\frac{1}{4}.$ (5)
@@
    $\textstyle 2LH\sigma^{2}\eta^{3}+\frac{\sigma^{2}}{M}\eta^{2}-\frac{D^{2}}{RH}=0.$
@@
    $\textstyle\gamma_{B}(\eta)=1+\frac{1}{H}\left(\frac{1}{4L\eta}-1\right),$
@@
    $\textstyle-\frac{4L^{2}D^{2}(H-1)}{R}+2L\sigma^{2}H\eta\left(\eta L(H-1)+1\right)^{2}+\frac{\sigma^{2}(H-1)}{MH}(\eta L(H-1)+1)^{2}=0.$
@@
The proof of the above proposition is straightforward and follows by writing the KKT conditions for the optimization problem in Equation˜5. A consequence of Proposition˜1 is that in the case of ill-tuning of the inner stepsize $\eta$, a large outer stepsize $\gamma$ can make up for it. For example, if $\sigma\rightarrow 0$ and $\eta LH\ll\mathcal{O}(1)$, we can make up for this by choosing $\gamma$ as $\frac{1}{\eta LH}$. Thus, we can interpret the outer learning rate $\gamma$ as having two dual roles. (a) It allows us to interpolate between minibatch SGD $(\gamma>1)$ and vanilla Local SGD $(\gamma=1)$, giving us the better of the two rates, and (b) it provides us some additional leeway in hyperparameter tuning by making up for ill-tuned inner learning rate $\eta$.
Our theory suggests that in the worst case, choices of $\gamma<1$ are not useful from an optimization perspective. We should either choose $\gamma=1$ or $\gamma>1$. This can be seen even on quadratic objectives, for example if $f(x)=\frac{x^{\top}Qx}{2}$ for some positive definite matrix $Q$, then a straightforward computation gives the expected iterate after $H$ local steps and $R$ communication rounds is $\mathbb{E}\left[x_{R}\right]=((1-\gamma)I+\gamma(I-\eta Q)^{H})x_{0}$. From this, it is clear that if $\eta$ is chosen such that $(I-\eta Q)^{H}$ has eigenvalues smaller than $1$, we should choose $\gamma\geq 1$. While if $(I-\eta Q)^{H}$ has any eigenvalues larger than $1$, we should just choose $\gamma=0$ (i.e. just don’t apply the algorithm at all). In other words, $\gamma$ can make up for a learning rate that is too small, but not a learning rate that is too large. This observation does not exclude that $\gamma<1$ can be useful from a generalization perspective, as noted for the case of a single client by [zhou21_lookahead], in the presence of data heterogeneity, as noted by [charles21_conver_accur_trade_offs_feder], or in the presence of specific stochastic gradient distributions (see Section˜3.4).
Proof sketch for Theorem˜1. We first start by expanding the update for the round iterate $x_{r+1}-x_{\ast}=x_{r+1}-x_{r}+x_{r}-x_{\ast}$ similar to [karimireddy19_scaff] to get,
@@
where $g_{r,h}$ is defined as in Equation˜1. [karimireddy19_scaff, jhunjhunwala23_fedex] control the inner product $\left\langle x_{r}-y_{r,h},g_{r,h}\right\rangle$ by either using smoothness or Young’s inequality; This would force us to bound the stray ${\left\lVert y_{r,h}-x_{r}\right\rVert}^{2}$ and take the local stepsize $\eta$ to be small in order to ensure convergence. Instead, we rely on bounding this quantity directly by viewing it as the regret in the online convex optimization sense with respect to the comparator $x_{r}$. Observe that the virtual sequence of averaged local iterates satisfies $y_{r,h+1}=y_{r,h}-\eta g_{r,h}$, and thus through standard regret analysis we have
$\textstyle\sum_{h=0}^{H-1}-\left\langle x_{r}-y_{r,h},g_{r,h}\right\rangle$ $=\frac{-{\left\lVert y_{r,h}-x_{r}\right\rVert}^{2}}{2\eta}+\frac{\eta}{2}\sum_{h=0}^{H-1}{\left\lVert g_{r,h}\right\rVert}^{2}.$ (6)
The negative terms ${\left\lVert y_{r,H}-x_{r}\right\rVert}^{2}$ in Equation˜6 turn out to be crucial in obtaining an analysis that works for all $\eta$ and not just small $\eta$. With this change and through carefully bounding the variance terms following [khaled19_tight_theor_local_sgd_ident_heter_data, woodworth20_is_local_sgd_better_than_minib_sgd], we obtain the guarantee of Theorem˜1. The full proof is provided in Section˜B.2.
Comparison with results on related algorithms. [malinovsky22_server_side_steps_sampl_without] analyze a closely related variant of the algorithm that uses federated random reshuffling [mishchenko21_proxim_feder_random_reshuf] as a base. This is a significantly different algorithm that doesn’t allow for an arbitrary number of local steps $H$ and depends on $f$ posessing finite-sum structure. Nevertheless, we can still specialize [malinovsky22_feder_random_reshuf_with_compr_varian_reduc, Theorem 2] approximately to our setting, by using $H$ as the number of data points in an epoch. In our notation, their convergence guarantee reads
$\textstyle\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{\eta\gamma HR}+\eta^{2}H^{2}\sigma^{2}\right),$
@@
$\textstyle\mathbb{E}\left[f\left(\frac{1}{R}\sum_{r=0}^{R-1}x_{r}\right)\right]-f(x_{\ast})\leq\mathcal{O}\left(\frac{L{\left\lVert x_{0}-x_{\ast}\right\rVert}^{2}}{R}\right),$
this rate is the same as gradient descent for $R$ steps (since the finite-sum structure means that per-epoch we approximate one step of gradient descent when $\eta$ is small). A similar rate is derived in [li24_power_extrap_feder_learn, li24_conver_fedpr_with_extrap_inexac_prox] if we have access to the proximal operator (i.e. we can do many local steps $H$ on a modified objective). [li24_power_extrap_feder_learn] in particular show that an outer learning rate greater than $1$ can be particularly useful for improving the convergence of FedProx [li18_feder_optim_heter_networ] in the heterogeneous setting when the smoothness constant varies significantly between different clients.